What is the molarity of a solution containing 33.5 g of NaCl dissolved in 500. mL of water?


1
Expert's answer
2022-05-16T17:03:04-0400

Solution:

The molar mass of NaCl is 58.44 g/mol

Therefore,

Moles of NaCl = (33.5 g NaCl) × (1 mol NaCl / 58.44 g NaCl) = 0.573 mol NaCl


Molarity = Moles of solute / Liters of solution

Assume that Liters of solution = Liters of water = (500 mL) × (1 L / 1000 mL) = 0.5 L

Therefore,

Molarity of NaCl solution = (0.573 mol) / (0.5 L) = 1.146 mol/L

Molarity of NaCl solution = 1.146 M


Answer: The molarity of NaCl solution is 1.146 M